# Toward a Team of AI-made Scientists for Scientific Discovery from Gene Expression Data

### v2

@@ Abstract
Machine learning has emerged as a powerful tool for scientific discovery, enabling researchers to extract meaningful insights from complex datasets. For instance, it has facilitated the identification of disease-predictive genes from gene expression data, significantly advancing healthcare. However, the traditional process for analyzing such datasets demands substantial human effort and expertise for the data selection, processing, and analysis. To address this challenge, we introduce a novel framework, a Team of AI-made Scientists (TAIS), designed to streamline the scientific discovery pipeline. TAIS comprises simulated roles, including a project manager, data engineer, and domain expert, each represented by a Large Language Model (LLM). These roles collaborate to replicate the tasks typically performed by data scientists, with a specific focus on identifying disease-predictive genes. Furthermore, we have curated a benchmark dataset to assess TAIS’s effectiveness in gene identification, demonstrating our system’s potential to significantly enhance the efficiency and scope of scientific exploration. Our findings represent a solid step towards automating scientific discovery through large language models.
@@ 1 Introduction
Leveraging Large Language Models (LLMs) (Liu et al., 2023) as agents (Zhou et al., 2023b), this paper proposes a Team of AI-made Scientists (TAIS) to automatically simulate researchers’ work. TAIS agents will execute tasks like dataset selection, preprocessing, confounder factor correction, condition prediction, and analysis to identify disease-predictive genes under various conditions.
Given the pioneering nature and complexity of our aim, We then establish a golden standard benchmark with datasets comprising 457 disease-condition pairs to assess our TAIS method’s performance. This benchmark involves manual selection, analysis, and processing of datasets, as well as writing and executing code to identify predictive genes for disease status under various conditions.
@@ 2.1 Pipelines to Identify Genes Predictive of Disease Status
The pipeline to analyze gene expression data with ML begins with the selection of datasets, such as public repositories like GEO (Clough and Barrett, 2016) or TCGA (Tomczak et al., 2015). The next step is data preprocessing. This involves cleaning the data, handling missing values (Abusamra, 2013), removing entries with an empty record, and eliminating genes that are not recorded (Khondoker, 2006). This step ensures the quality and reliability of the data before analysis. Then, one can perform a basic regression analysis to identify the genes that are predictive of the disease status (Ghosh and Chinnaiyan, 2005; Wu et al., 2009). Lasso (Tibshirani, 1996) is often chosen as the model due to its ability to identify a sparse set of genes, which helps pinpoint a relatively small number of genes from thousands of possible predictors. Additional steps to correct potential confounding factors (Leek et al., 2010; Bruning et al., 2016) is usually needed for a more reliable identification of the genes, especially in the scenario where the distribution of gene expressions varies across subgroups within the data due to different background distributions, rather than the disease itself (Yu et al., 2006). Recent studies extend the discoveries pipeline to incorporate more conditions in the analysis (Yang et al., 2023b), including additional covariates in the regression model, such as patient demographics, and other possible co-occurrence of diseases (Kyalwazi et al., 2023). The inclusion of these conditions allows for the identification of gene expression patterns that are not only associated with the disease status but also modulated by these conditions. This nuanced analysis supports the development of more personalized treatment strategies by identifying how different conditions affect gene-disease relationships (Rosenquist et al., 2023).
@@ 2.2 Task Solving via LLMs as Agents
LLMs have illustrated the potential to achieve human-level intelligence (Wang et al., 2023b; OpenAI, 2023; Touvron et al., 2023a, b). Building upon this capability, subsequent research explores and enhances LLMs’ ability to solve complex tasks (Wang et al., 2022a, c; Hao et al., 2023; Yao et al., 2022). For example, CoT (Wei et al., 2022) undertake tasks through goal decomposition (Zheng et al., 2023; Feng et al., 2023; Wang et al., 2022a; Ning et al., 2023) and tackles each sub-goal sequentially. Given this ability, it’s expected that the collaboration of multiple agents could further enhance the problem-solving capacities (Wang et al., 2023c; Talebirad and Nadiri, 2023; Du et al., 2023; Wang et al., 2023a), with each agent performing specific actions with distinct expertise (Yang et al., 2023a; Talebirad and Nadiri, 2023; Dong et al., 2023). For example, MetaGPT (Hong et al., 2023) promotes collaboration among various roles of AI agents (e.g. product manager, engineer, etc) by SOPs in software development. Qian et al. (2023); Dong et al. (2023) has also demonstrated role-play in software development and code generation. Other works focus on sociological phenomena (Shapiro et al., 2023; Sumers et al., 2023; Zhou et al., 2023a; Wang et al., 2023c; Li et al., 2023). Researchers have also been trying to incorporate LLMs into scientific discovery such as in chemistry (Bran et al., 2023; Guo et al., 2023), biotechnology (Madani et al., 2023) and medical (Singhal et al., 2023; Yang et al., 2023c) areas by training or finetuning LLMs on domain-specific data. In contrast to these works, We leverage available LLMs out-of-box without any training and assemble a team of AI scientists tasked with the planning, analysis, and coding to solve scientific problems in genomics.
@@ 3 Method
3.1 Overview: the Team of AI-made Scientists
Our Team of AI-made Scientists (TAIS) consists of five roles: Project Manager, Data Engineer, Domain Expert, Statistician, and Code Reviewer, each contributing distinct expertise to solve scientific problems.
Figure 2: The collaboration between Data Engineer and Code Reviewer.
Figure 2: The collaboration between Data Engineer and Code Reviewer.
Figure 3: The collaboration between Data Engineer and Domain Expert.
For a detailed description of these roles, please refer to Appendix A. The overall collaboration scheme is illustrated in Figure 1, where three AI agents (Data Engineer, Domain Expert, and Code Reviewer) are tasked with the data preprocessing procedure, and two AI agents (Statistician and Code Reviewer) with the regression analysis part.
@@ 3.2 Collaboration among AI Scientists
We highlight two critical collaborations among AI scientists in TAIS. The first, termed program-and-review, involves the interaction between the Code Reviewer and either a Statistician or Data Engineer. Throughout the coding process, code from the Data Engineer or Statistician is reviewed by the Code Reviewer agent to ensure quality and adherence to instructions, as shown in Figure 3. This process continues until the code meets the review standards or reaches the maximum number of review rounds. The Code Reviewer evaluates the code based on its executability, error-free status, and compliance with the given instructions. Successful review allows the code to proceed; otherwise, detailed feedback is provided to the coder (e.g. Data Engineer and Statistician) for revision.
The second collaboration, Consultative Coding, involves a Data Engineer working alongside a Domain Expert who offers domain-specific knowledge. For example, in a project focusing on extracting clinical information from breast cancer data, the Data Engineer managed basic data conversions but faced challenges with acronyms. Consultation with a Domain Expert introduced a method to parse complex medical terms into the expected data type, addressing issues like identifying acronyms of the Triple Negative Breast Cancer (TNBC) subtypes. This collaborative effort led to the successful completion of missing data, underscoring the importance of teamwork and expert insight in enhancing data preprocessing.
3.3 Regression Analysis on Gene Expression Data
Analyzing gene expression data to identify significant gene factors, considering confounding variables, involves a comprehensive statistical approach due to the data’s high-dimensionality and heterogeneity.
@@ 3.3 Regression Analysis on Gene Expression Data
Further, to incorporate additional conditions such as age and gender into the analysis, the residualization approach is employed to account for the effect of the condition. For cases lacking direct condition data, a two-step regression strategy is adopted. This involves using common genes between datasets to estimate missing conditions, thus enabling the integration of trait and condition data for comprehensive analysis. Detailed methodologies and statistical models are discussed in Appendix B.
@@ 4 Benchmark Creation
A computational biology researcher identified a list of key biomedical entities related to genetics research or public health, resulting in 65 traits classified into 9 categories. These traits were paired with either another trait or demographic attributes like “age” or “gender”, generating 4556 possible pairs. These pairs were designed to pose questions in the format: “What are the significant genes related to the trait when considering the influence of the condition?” Then, we used a set of inclusion and exclusion criteria (Appendix C) and ranked the trait-condition association of pairs based on the Jaccard similarity of related genes from the NCBI Gene database, to identify 457 pairs that are of most scientific interest, which form our benchmark’s question set. Details on these pairs are available in Appendix D.
@@
Gene expression and clinical data were obtained from the Gene Expression Omnibus (GEO) (Clough and Barrett, 2016) and the Cancer Genome Atlas (TCGA) (Tomczak et al., 2015) through the Xena platform (Goldman et al., 2020). Gene symbols related to traits were sourced from the NCBI Gene database (Brown et al., 2015). For more information on these data sources, see Appendix E.
@@ 5.1 Experiment Setting
To evaluate the performance of TAIS in identifying trait-related genes, we ran our system with the raw datasets as input and compared the identified genes with those from the gold standard, using metrics such as precision, recall, and Jaccard index. We started with an end-to-end evaluation on the whole data analysis process, and then we separately analyzed the preprocessing and regression abilities of our system.
@@ 5.2 Main Results
We first report the main results in Section 5.2.1. Then, to further analyze the performance of TAIS on each sub-problem, we provide ablation analysis on regression analysis in Section 5.2.2 and data preprocessing 5.2.3.
@@ 5.2.1 Performance of TAIS System
Table 1 demonstrates the performance of TAIS on our benchmark. We respectively display the performance on single-step and two-step regression problems and the overall performance. We achieve 67.74% Success Rate, 9.20% Precision and 10.40% Recall on the single-step problems. Two-step regression is relatively more difficult, which is reflected on the performance, i.e. 42.28% Success Rate, 6.49 % Precision and 6.78% Recall. Given the difficulty of the problem, the results we achieve are impressive. A qualitative analysis in Appendix F on a specific question reveals that genes pinpointed by TAIS align with findings from biomedical studies. We also experiment with other SOTA methods, e.g. MetaGPT and AutoGPT, on our benchmarks. We find that these methods are not capable of generating bug-free code on the data preprocessing. Thus, we don’t report their results here.
@@
Success Rate (%) Precision (%) Recall (%) Jaccard (%)
Single-step 67.74 9.20 10.40 6.73
Two-step 42.28 6.49 6.78 4.18
Overall 45.73 6.86 7.28 4.53
@@ 5.2.2 Performance of Regression
To validate the ability of TAIS to perform the regression analysis to identify genes that are predictive of disease status, we replace the TCGA and GEO datasets preprocessed by our Data Engineer with the gold standard ones. Using these preprocessed data as input, we evaluated our system on the regression analysis sub-task.
@@
Budget (# reviews) Success Rate (%) Precision (%) Recall (%) Jaccard (%)
Single-step 0 0.00 0.00 0.00 0.00
1 95.16 34.60 35.99 23.90
2 95.16 34.60 35.99 23.90
Two-step 0 83.54 2.08 2.06 1.19
1 5.57 0.14 0.14 0.08
2 83.54 1.81 1.83 1.09
Overall 0 72.21 1.80 1.78 1.03
1 17.72 4.82 5.00 3.31
2 85.12 6.26 6.46 4.19
Performance Table 2 demonstrates results of TAIS on regression sub-task only. As per the table, our TAIS shows a much higher Success Rate on the regression sub-task with the preprocessed data from golden standards compared with results in Table 1. Among different budgets, we achieve the best results of 85.12% Success Rate, 6.26% Precision and 6.46% Recall on overall regression problems.
Comparison with Baselines. We adopt the recent methods on LLM agents and Code Generation and conduct experiments on our benchmark. We experiment with two settings: 1. Instruct with a general task description 2. Instruct with a detailed task description, the same as ours. We experimented with MetaGPT and AutoGPT and found that despite the feedback mechanism in respective method, the generated code leads to various persistent errors in the pipeline (such as NaN during regression) and no testable output can be generated.
Different maximum number of reviews. As discussed in Section 3.2, a Code Reviewer is recruited for any task that involves code generation. During the process, multiple rounds of program-and-review are performed. One expects that with more rounds (i.e. more budgets), the generated code will be of higher quality and with fewer bugs. We perform an analysis on the budgets and demonstrate the results under different budges in Table 2. As can be seen, with more budgets, a better Success Rate and Precision-Recall can be achieved.
@@
Budget (# reviews) Success Rate (%) Precision (%) Recall (%) Jaccard (%)
Single-step 0 25.80 3.06 3.03 2.37
1 32.25 2.98 3.04 2.34
2 95.20 34.60 36.00 23.90
Two-step 0 49.87 9.75 9.81 7.76
1 59.49 11.27 11.37 8.91
2 83.50 1.81 1.83 1.10
Overall 0 46.60 8.84 8.89 7.02
1 55.80 10.15 10.21 8.02
2 85.10 6.30 6.50 4.20

### v3

@@ Abstract
Machine learning has emerged as a powerful tool for scientific discovery, enabling researchers to extract meaningful insights from complex datasets. For instance, it has facilitated the identification of disease-predictive genes from gene expression data, thereby improving risk stratification, early diagnosis, and treatment selection. However, the traditional process for analyzing such datasets demands substantial human effort and expertise for the data selection, processing, and analysis. To address this challenge, we introduce a novel framework, a Team of AI-made Scientists (TAIS), designed to streamline the scientific discovery pipeline. TAIS comprises simulated roles, including a project manager, data engineer, and domain expert, each represented by a Large Language Model (LLM). These roles collaborate to replicate the tasks typically performed by data scientists, with a specific focus on identifying disease-predictive genes. Furthermore, we have curated a benchmark dataset to assess TAIS’s effectiveness in gene identification, demonstrating our system’s potential to significantly enhance the efficiency and scope of scientific exploration. Our findings represent a solid step towards automating scientific discovery through large language models.  [^1] Code for a more recent version of our system is available at https://github.com/Liu-Hy/GenoMAS.
@@ 1 Introduction
Leveraging Large Language Models (LLMs) (Patil and Gudivada, 2024) as agents (Guo et al., 2024), this paper proposes a Team of AI-made Scientists (TAIS) to automatically simulate researchers’ work. TAIS agents will execute tasks like dataset selection, preprocessing, confounder factor correction, condition prediction, and analysis to identify disease-predictive genes under various conditions.
Given the pioneering nature and complexity of our aim, We then establish a gold standard benchmark with datasets comprising 457 disease-condition pairs to assess our TAIS method’s performance. This benchmark involves manual selection, analysis, and processing of datasets, as well as writing and executing code to identify predictive genes for disease status under various conditions.
@@ 2.1 Pipelines to Identify Genes Predictive of Disease Status
The pipeline to analyze gene expression data with ML begins with dataset selection (e.g., GEO (Clough and Barrett, 2016) or TCGA (Tomczak et al., 2015)) and preprocessing: cleaning, handling missing values (Abusamra, 2013), removing empty records, and excluding unrecorded genes (Khondoker, 2006). One then fits regression models to identify genes predictive of disease status (Ghosh and Chinnaiyan, 2005; Wu et al., 2009), often using Lasso for sparsity (Tibshirani, 1996). To reduce bias, pipelines correct confounding/batch effects (Leek et al., 2010; Bruning et al., 2016; Yu et al., 2006). Later work further integrates covariates (demographics, comorbidities) to enable precision analyses (Yang et al., 2023b; Kyalwazi et al., 2023; Rosenquist et al., 2023). More recently, several studies refined these pipelines along complementary axes. Comparative work assessed how preprocessing choices impact cross-study generalization for transcriptomic prediction, underscoring the importance of harmonized workflows (Mize et al., 2024). Methodologically, deep learning and graph-based models improved disease–gene association discovery from omics graphs and sequence-derived representations (Saadat and Fellay, 2024). At the interface of genomics and pathology, recent reviews synthesize advances in deep learning that connect molecular profiles with histopathology, informing pipeline design for robust biomarker discovery (Unger and Kather, 2024).
@@ 2.2 Task Solving via LLMs as Agents
LLMs show strong general capabilities (Wang et al., 2023b; OpenAI, 2023; Touvron et al., 2023a, b). Work on reasoning and acting (Wang et al., 2022a, c; Hao et al., 2023; Yao et al., 2022) popularized CoT (Wei et al., 2022) with goal decomposition (Zheng et al., 2023; Feng et al., 2023; Wang et al., 2022a; Ning et al., 2023). Multi-agent systems further amplify problem solving (Wang et al., 2023c; Talebirad and Nadiri, 2023; Du et al., 2023; Wang et al., 2023a; Yang et al., 2023a; Dong et al., 2023), and role-based frameworks like MetaGPT operationalize collaboration (Hong et al., 2023; Qian et al., 2023). As evaluation has matured in 2024, new benchmarks probe complementary dimensions: dynamic multi-agent competence (LLMArena) (Chen et al., 2024), cooperation/competition (BattleAgentBench) (Wang et al., 2024), safety risks in interactive settings (Agent-SafetyBench) (Zhang et al., 2024), and progress tracking via modular tasks and metrics (AgentQuest) (Gioacchini et al., 2024). Beyond enabling software tasks, recent efforts target end-to-end science: the AI Scientist proposes automated idea generation, experimentation, and paper writing (Lu et al., 2024). Complementary, forecasting pipelines use evolving knowledge graphs to predict emergent impactful directions (Gu and Krenn, 2024). In contrast to domain-specific finetuning in chemistry/biotech/medicine (Bran et al., 2023; Guo et al., 2023; Richard et al., 2024; Tang et al., 2024), we leverage off-the-shelf LLMs and coordinate a team for genomics.
@@ 3 Method
3.1 System overview: a lightweight role-driven team
TAIS organizes a small set of specialized agents into a two-stage pipeline for gene expression analysis: data preparation followed by regression-based association testing. The team comprises five roles with minimal but complementary responsibilities: Project Manager (coordinator) parses the user query (trait and optional condition), scopes required datasets (e.g., TCGA, GEO), and schedules two sequential stages with checkpoints. Data Engineer implements dataset-specific preprocessing code. Statistician runs regression to identify trait-associated genes while accounting for confounding. Domain Expert acts as a biomedical consultant for decisions that hinge on domain knowledge (e.g., cohort inclusion, clinical variable parsing, gene symbol normalization). Code Reviewer audits generated code for executability and instruction conformance.
Execution is intentionally simple. The Project Manager issues stage descriptions and acceptance criteria; the Data Engineer and Statistician write short code segments, execute them, and submit outputs and logs. The Code Reviewer provides bounded feedback (a small, fixed number of review rounds) when code fails or drifts from instructions. The Domain Expert is queried only at decision points requiring biomedical judgment (e.g., mapping histology strings to labels, platform-specific gene identifier handling). This lightweight design avoids heavy orchestration machinery while still enforcing basic quality control.
Figure 2: Write–run–audit loop between the Data Engineer and Code Reviewer.
Figure 3: Consultation between the Data Engineer and Domain Expert on biomedical decisions.
For brief role descriptions, see Appendix A. An overview schematic is shown in Figure 1. In TAIS, three agents (Data Engineer, Domain Expert, Code Reviewer) collaborate on preprocessing, while the Statistician and Code Reviewer handle regression.
@@ 3.2 Collaboration among AI Scientists
We employ two interaction patterns:
Write–Run–Audit (program-and-review). For any step that generates code, the authoring agent (Data Engineer or Statistician) executes the code and submits the snippet, stdout/stderr, and step instruction to the Code Reviewer (Figure 3). The reviewer checks two things: (i) whether the code runs without errors and (ii) whether it follows the instruction and acceptance criteria. If rejected, the author revises and re-runs. The loop is bounded by a small review budget to limit latency and overfitting to feedback.
Consultative Coding. When a step depends on biomedical knowledge (e.g., extracting clinical labels from free-text metadata, merging gene identifiers across platforms), the Data Engineer requests guidance from the Domain Expert (Figure 3). The expert returns concise, actionable advice or pseudo-code that the engineer turns into executable code. If execution fails, the consultation can repeat within the same step until the review budget is exhausted.
End-to-end flow.
In the end-to-end analysis flow, given a query (trait and optional condition), the Project Manager (i) locates candidate cohorts, (ii) chooses stage order and checkpoints, and (iii) starts preprocessing. The Data Engineer performs file parsing, sample filtering, clinical feature extraction, gene symbol normalization, and normalization of expression values. The outputs are matrix-shaped tables with aligned sample identifiers and optional condition columns. After the audit passes, the Statistician selects a regression recipe (Section 3.3) based on basic diagnostics and produces a ranked list of genes with effect estimates. All steps are auditable and rerunnable with the same inputs and random seeds.
3.3 Regression on gene expression with basic confounding control
Analyzing gene expression data to identify significant gene factors, considering confounding variables, involves a comprehensive statistical approach due to the data’s high-dimensionality and heterogeneity. We summarize the statistical backbone used by the Statistician.
@@ 3.3 Regression on gene expression with basic confounding control
Further, to incorporate additional conditions such as age and gender into the analysis, the residualization approach is employed to account for the effect of the condition. For cases lacking direct condition data, a two-step regression strategy is adopted. This involves using common genes between datasets to estimate missing conditions, thus enabling the integration of trait and condition data for comprehensive analysis.
@@ 4 Benchmark Creation
A computational biology researcher identified a list of key biomedical entities related to genetics research or public health, resulting in 65 traits classified into 9 categories. These traits were paired with either another trait or demographic attributes like “age” or “gender”, generating 4556 possible pairs. These pairs were designed to pose questions in the format: “What are the significant genes related to the trait when considering the influence of the condition?” Then, we used a set of inclusion and exclusion criteria (Appendix B) and ranked the trait-condition association of pairs based on the Jaccard similarity of related genes from the NCBI Gene database, to identify 457 pairs that are of most scientific interest, which form our benchmark’s question set. Details on these pairs are available in Appendix C.
@@
Gene expression and clinical data were obtained from the Gene Expression Omnibus (GEO) (Clough and Barrett, 2016) and the Cancer Genome Atlas (TCGA) (Tomczak et al., 2015) through the Xena platform (Goldman et al., 2020). Gene symbols related to traits were sourced from the NCBI Gene database (Brown et al., 2015). For more information on these data sources, see Appendix D.
To enhance the evaluation of our TAIS method and facilitate future research utilizing large language models in genetic data exploration, we introduce a benchmark dataset comprised of 457 meticulously formulated questions, alongside a gold standard for resolving each, which encompasses genetic datasets sourced from public repositories, preprocessing and regression analysis code, and corresponding results. Our benchmark is named as the Genetic Question Exploration (GenQEX) dataset. This section delineates the benchmark’s development process.
Question Generation
A computational biology researcher curated a list of significant biomedical entities, encompassing human traits or diseases pivotal to genetics research or public health. After a thorough manual curation, a diverse list of 67 traits across 9 primary categories was compiled. Subsequently, each trait was paired with a condition—either another trait from the list or one of the demographic attributes “age” or “gender”—resulting in 4556 potential pairs. These pairs were designed to pose questions in the format: “What are the significant genes related to the trait when considering the influence of the condition?” To sift through these pairs for scientifically relevant queries, we employed specific inclusion and exclusion criteria. The pairs were then ranked based on the Jaccard similarity between the related genes of the trait and the condition, derived from the NCBI Gene database, to identify pairs where the trait is implicated with the condition. This process yielded 457 pairs, which collectively form our benchmark’s question set. For a comprehensive list of these pairs please see Appendix C.
Input Dataset
To address the formulated research questions, gene expression and corresponding clinical data were procured from renowned public databases: The Cancer Genome Atlas (TCGA) via the Xena platform, and the Gene Expression Omnibus (GEO). Additionally, domain knowledge regarding gene symbols associated with traits was sourced from the NCBI Gene database. Please refer to Appendix D for an introduction about these data sources.
@@ 5.1 Experiment Setting
We evaluate TAIS on our benchmark of gene-trait association tasks using five metrics: Success Rate (SR), Precision, Recall, F1, and Jaccard. We consider three complementary protocols to isolate contributions of each stage: (1) end-to-end: TAIS performs both preprocessing and regression; (2) regression-only: we replace the inputs with gold-standard preprocessed data to assess the Statistician in isolation; (3) preprocessing-only: we feed data preprocessed by TAIS into a gold-standard regression script to assess the Data Engineer in isolation. For stages involving code generation, we vary the review budget (maximum number of write-run-audit rounds) as defined in Section 3.2.
@@ 5.2 Main Results
We first present the end-to-end results in Section 5.2.1. We then present evaluation of the regression-only setting in Section 5.2.2 and the preprocessing-only setting in Section 5.2.3.
@@ 5.2.1 Performance of TAIS System
Table 1 summarizes end-to-end performance. Overall, TAIS attains SR 69.08%, Precision 33.91%, Recall 31.70%, F1 30.27%, and Jaccard 21.13%. The single-step setting is substantially easier than the two-step setting (F1 45.05% vs. 19.21%), reflecting the difficulty of estimating or integrating missing conditions across cohorts. These results are consistent with our design choices: a lightweight team with bounded review budgets yields reasonable performance, but struggles most when preprocessing must infer conditions.
@@
Success Rate (%) Precision (%) Recall (%) F1 (%) Jaccard (%)
Single-step 71.27 48.35 43.84 45.05 30.15
Two-step 67.43 23.01 22.55 19.21 14.33
Overall 69.08 33.91 31.70 30.27 21.13
@@ 5.2.2 Performance of Regression
To evaluate the Statistician, we use gold-standard preprocessed datasets as inputs and vary the review budget for program-and-review. Results in Table 2 show that code review markedly improves performance: overall F1 increases from 55.95% (0 reviews) to 80.12% (1 review) and 89.30% (2 reviews). Gains are consistent across single-step (F1 58.61% to 94.74%) and two-step (53.77% to 85.20%), indicating that most of the statistical errors are correctable through bounded iteration. The Statistician’s logic is thus strong when provided with clean inputs.
@@
Budget (# reviews) Success Rate (%) Precision (%) Recall (%) F1 (%) Jaccard (%)
Single-step 0 62.56 58.85 58.45 58.61 43.13
1 87.97 87.02 84.84 85.62 77.54
2 97.45 95.88 93.71 94.74 89.38
Two-step 0 58.80 56.22 51.88 53.77 38.59
1 83.87 78.31 75.19 74.92 63.24
2 91.33 85.88 84.44 85.20 76.11
Overall 0 60.42 57.25 54.71 55.95 41.68
1 85.63 81.05 79.34 80.12 69.39
2 93.96 90.18 88.43 89.30 81.83
The sharp jump from 0 to 1 review highlights the value of the audit loop for catching implementation drift and minor numerical issues; the second review yields diminishing but still notable returns. Since inputs are fixed and clean, the remaining gap to end-to-end performance mainly stems from preprocessing quality.
@@
Budget (# reviews) Success Rate (%) Precision (%) Recall (%) F1 (%) Jaccard (%)
Single-step 0 36.62 23.72 22.81 21.93 16.97
1 68.98 40.95 39.17 40.03 25.63
2 78.95 50.36 46.89 47.17 32.07
Two-step 0 33.48 10.45 9.62 8.07 6.19
1 59.54 22.09 21.95 18.04 14.28
2 70.30 26.24 26.76 22.33 16.66
Overall 0 34.83 16.16 15.29 14.41 10.83
1 63.59 30.20 29.35 27.50 19.16
2 74.02 36.61 35.42 33.01 23.29